a) What is the premise of ABC Analysis? What are the three steps in making ABC...

  1. a) What is the premise of ABC Analysis? What are the three steps in making ABC inventory analysis?

  2. b) State and describe with examples the three types of costs associated with inventory

  3. c) The annual demand of a Bottled Water Distribution company is 25,000 boxes, the ordering cost ¢40 per order, the carrying cost 15%, and the unit cost ¢15 per box. The order quantity is 500 units. Calculate the following;

    i). Annual ordering cost ii). Annual carrying cost iii). Total annual cost

  4. + iv). Total number of orders per year
Answer #1

a)

ABC analysis is an inventory categorization technique in which items are classified into A, B, and C class items based on their total consumption or revenue values. A class items are the limited high-value items that contribute to generally around 75%- 80% of the total consumption or revenue values. B class items are mid values items that contribute to 15%-20% of the total consumption or revenue values. C class items are the least value items that contribute to 5%-10% of the total consumption or revenue values. Generally, A class items are lower in numbers than B and C class items.

The steps in making ABC inventory analysis are:

1. Calculate the consumption or revenue values for each item and their percentage of total consumption or revenue value.

2. Arrange the items in the descending order of the percentage of total consumption or revenue value. The cumulative percentage of total consumption or revenue value is then calculated for each item.

3. Classify the items as A, B, and C. A items contribute to top 80%, B items 80%-95%, and C items 95%-100% of the cumulative percentage of total consumption or revenue value.

b)

The three types of cost associated with inventory are:

1. Purchase cost: Purchase cost is the total cost of inventory items over a period of time.

For example, if the annual demand for an item is 5000, and the per-unit price of the item is $5:

Annual Purchase cost = 5000 x 5 = $ 25000

2. Carrying cost: It is the cost of holding or carrying inventory over a period of time.

For example, if the average inventory in a year is 200 units, and the per-unit carrying cost for one year is $2:

Annual Holding cost = 200 x 2 = $ 400

3. Ordering cost: It is the cost of ordering the inventory items over a period of time.

For example, if the number of orders in a year is 10, and the ordering cost per order is $5:

Annual Ordering cost = 10 x 5 = $50.

c.

Annual Demand = 25000 boxes

Ordering cost = $ 40 per order

Unit cost = $15

Carrying cost = 15% of Unit cost = 0.15 x 15 = $ 2.25

Order Quantity = 500 units

Number of orders in a year = Annual Demand / Order Quantity = 25000 / 500 = 50

i)

Annual Ordering cost = Number of orders in a year x Ordering cost per order = 50 x 40 = $ 2000

ii)

Average Inventory = Order Quantity / 2

Annual Carrying cost = Average Inventory x Carrying cost per unit = (500 / 2) x 2.25 = $ 562.5

iii)

Total Annual Cost = Annual Purchasing cost + Annual Ordering Cost + Annual Carrying cost

Annual Purchasing cost = Annual Demand x Unit cost = 25000 x 15 = $ 375000

So,

Total Annual Cost = 375000 + 2000 + 562.5 = $ 377562.5

iv)

Total number of orders in a year = Annual Demand / Order Quantity = 25000 / 500 = 50